# Yichen Yu1,2       Huan-Song Xu311footnotemark: 1

### v1

@@ Abstract
1 INTRODUCTION
@@
In typical scenarios, visually impaired runners must rely on a human guide, often tied by a short rope or connected by a series of verbal cues. Synchronization between the guide and the visually impaired runner is essential but often difficult to maintain. Verbal guidance demands constant attention and clear environmental acoustics, which may not always be feasible in outdoor environments. Moreover, the use of physical tethers can restrict natural stride, limit arm swing, and introduce discomfort, ultimately disrupting the running rhythm and reducing user agency. These traditional methods also reinforce a unidirectional dependency model, where the guide dictates the pace and direction, leaving the visually impaired individual with limited control over their experience Xu et al. (2017); Lind et al. (2020).
@@
Emerging research in motor coordination and haptic feedback systems has demonstrated potential in reducing these barriers. Studies have demonstrated that rhythmic tactile cues, when delivered with precision, can reinforce movement regularity and support spontaneous synchronization between individuals. This phenomenon, known as entrainment, reflects the human body’s innate ability to synchronize with external rhythmic stimuli and has been observed across a range of social and locomotor activities Islam and Lim (2022); Schroeder et al. (2025); Georgiou et al. (2020).
@@
To leverage this potential, we present RunPacer, an innovative smartwatch-based system intended to promote real-time, bidirectional synchronization between a guide and a visually impaired runner. Rather than relying on verbal instructions or physical tethers, RunPacer employs shared vibrotactile feedback, delivering synchronized step-based pulses to both participants. This approach offers a seamless, intuitive, and cognitively light alternative to traditional methods, empowering visually impaired enabling users to participate in running experiences that are not only safer and more comfortable but also more autonomous, social, and enjoyable Baldi et al. (2020); Shull and Damian (2015); Ballardini et al. (2020).
2 RELATED WORK
@@
Researchers have long explored the use of haptic and vibrotactile feedback to guide motion in rehabilitation, exercise, and assistive technologies, aiming to support users with varying motor or sensory limitations. A considerable body of work has demonstrated that haptic cues can be pivotal in motor learning, spatial orientation, and rhythm synchronization. For example, Xu et al. developed a wearable vibrotactile system that improves postural control through directional feedback, showing encouraging outcomes in both static balance tasks and dynamic movements Xu et al. (2017). Similarly, Kalff et al. found that gait-synchronized vibrotactile feedback enhances mobility results in subjects with lower-limb amputations, demonstrating how real-time tactile feedback can facilitate neuromotor rehabilitation and entrainment Kalff et al. (2024).
@@
Beyond medical rehabilitation, haptic feedback has been increasingly integrated into locomotion assistance systems. In the domain of gait training, De Angelis et al. performed an exhaustive analysis and concluded that vibrotactile stimulation not only supports postural stability but also encourages proactive participation by providing non-intrusive cues that users can internalize De Angelis et al. (2021). Mahmud et al. extended this notion to immersive environments by incorporating vibrotactile signals into virtual reality standing balance training programs, showing that tactile stimuli could effectively substitute or augment visual cues Mahmud et al. (2022a, b).
@@
Wearable haptic systems have also been used to facilitate socially coordinated behaviors. Lisini et al., for instance, proposed a wearable solution for remote social walking, allowing pairs of persons to preserve synchronized step rhythms even while being geographically separated Baldi et al. (2020). Their work illustrates the potential of vibrotactile entrainment in interpersonal contexts, suggesting broader applications beyond solo rehabilitation, including shared physical activities like dancing, walking, and running.
@@
However, most prior work focuses on navigation, balance, or motor control in constrained indoor settings, often with the goal of compensating for deficits or preventing falls. Assistive devices designed for the visually impaired, such as Wayband and other commercial systems, have largely targeted directional guidance, providing users with cues for pathfinding or obstacle avoidance through vibrotactile feedback embedded in wearables like shoes or belts. While effective for spatial navigation, these systems are not optimized for interpersonal movement synchronization, particularly in high-tempo and coordination-intensive scenarios like running with a guide Lind et al. (2020); Islam (2018).
@@
Moreover, despite the exhibited promise of entrainment theory in social and sensorimotor contexts, few systems have explicitly operationalized this concept into shared, bidirectional pacing tools for co-locomotion. The majority of existing solutions emphasize asymmetrical interaction models, where one user leads and the other passively follows. This dynamic limits autonomy and co-agency, especially for visually impaired individuals who may wish to take a more active role in the running experience Islam and Lim (2022).
@@
Our work bridges multiple strands of this literature—rehabilitation, wearable haptics, social entrainment, and assistive technology—to fill a void: the lack of tools that support symmetric, co-tempo running experiences for visually impaired individuals. RunPacer introduces a new category of assistive systems that use synchronized haptic feedback not for navigation, but for enabling mutual rhythm alignment. This repositioning shifts the emphasis from guidance to collaboration, reframing the role of the assistive device as a facilitator of shared agency and embodied communication Shull and Damian (2015); Ballardini et al. (2020).
3 SYSTEM DESIGN
@@
Figure 1: Overview of the RunPacer system, composed of three interconnected modules that operate in tandem to support synchronized co-running.
@@
In either case, a synchronized pulse stream is transmitted to both runners via a low-latency Bluetooth channel. The synchronization system is engineered to operate with under 100 milliseconds of delay, which is essential for preserving the perceptual integrity of the rhythm and maintaining entrainment Schroeder et al. (2025). The cadence engine is modular and platform-agnostic, supporting integration across smartwatch ecosystems using standardized communication APIs.
@@
Vibrotactile Feedback Delivery, transforms the rhythmic signals into distinct, synchronized pulses using the built-in haptic motors of each smartwatch Georgiou et al. (2020). These pulses are short, crisp, and calibrated to provide perceptual clarity without becoming intrusive or fatiguing.
The key design parameters include pulse duration, intensity, and waveform configuration. The pulse duration is optimized between 100–200 milliseconds to ensure clarity during motion. Pulse intensity is adjustable to accommodate ambient noise levels, individual preferences, and individual skin sensitivity Xu et al. (2017). The pulse waveform is configured to deliver single-tap pulses, which were favored over longer or multi-tap patterns in our preliminary user testing.
@@ 7 CONCLUSION
The contribution of RunPacer lies not only in its specific implementation, but also in the broader design implications it offers for inclusive movement interfaces. It reframes co-running as a bidirectional, rhythmic, and social interaction, leveraging entrainment principles and vibrotactile cues to support mutual adaptation rather than passive following. This reframing has possible implementations beyond running, including co-walking, synchronized rehabilitation, or embodied collaborative learning Ballardini et al. (2020).
@@
Acknowledgments
We thank Feng Chia University and Zhejiang University for their support in the early stages of this research. We also gratefully acknowledge Apple Inc. for providing technical resources and inspiration through the iOS Mobile Application Innovation Competition. Special thanks to Professor Ming-Yen Lin for his continuous mentorship and encouragement throughout the development of this project.

### v2

# RunPacer: A Smartwatch-Based Vibrotactile Feedback System for Symmetric Co-Running by Visually Impaired Individuals and Guides

@@ Abstract.
Accessibility, Assistive Technology, Vibrotactile Feedback, Wearable Computing, Haptic Interaction, Co-running, Smartwatch
journalyear: 2025copyright: rightsretainedconference: The 27th International ACM SIGACCESS Conference on Computers and Accessibility; October 26–29, 2025; Denver, CO, USAbooktitle: The 27th International ACM SIGACCESS Conference on Computers and Accessibility (ASSETS ’25), October 26–29, 2025, Denver, CO, USAdoi: 10.1145/3663547.3759738isbn: 979-8-4007-0676-9/2025/10ccs: Human-centered computing Accessibility technologiesccs: Human-centered computing Ubiquitous and mobile computing systems and toolsccs: Hardware Haptic devices
1. Introduction
@@
In typical scenarios, visually impaired runners must rely on a human guide, often tied by a short rope or connected by a series of verbal cues. Synchronization between the guide and the visually impaired runner is essential but often difficult to maintain. Verbal guidance demands constant attention and clear environmental acoustics, which may not always be feasible in outdoor environments. Moreover, the use of physical tethers can restrict natural stride, limit arm swing, and introduce discomfort, ultimately disrupting the running rhythm and reducing user agency. These traditional methods also reinforce a unidirectional dependency model, where the guide dictates the pace and direction, leaving the visually impaired individual with limited control over their experience (Xu et al., 2017; Lind et al., 2020).
@@
Emerging research in motor coordination and haptic feedback systems has demonstrated potential in reducing these barriers. Studies have demonstrated that rhythmic tactile cues, when delivered with precision, can reinforce movement regularity and support spontaneous synchronization between individuals. This phenomenon, known as entrainment, reflects the human body’s innate ability to synchronize with external rhythmic stimuli and has been observed across a range of social and locomotor activities (Islam and Lim, 2022; Schroeder et al., 2025; Georgiou et al., 2020).
@@
To leverage this potential, we present RunPacer, an innovative smartwatch-based system intended to promote real-time, bidirectional synchronization between a guide and a visually impaired runner. Rather than relying on verbal instructions or physical tethers, RunPacer employs shared vibrotactile feedback, delivering synchronized step-based pulses to both participants. This approach offers a seamless, intuitive, and cognitively light alternative to traditional methods, empowering visually impaired enabling users to participate in running experiences that are not only safer and more comfortable but also more autonomous, social, and enjoyable (Baldi et al., 2020; Shull and Damian, 2015; Ballardini et al., 2020).
2. Related Work
@@
Researchers have long explored the use of haptic and vibrotactile feedback to guide motion in rehabilitation, exercise, and assistive technologies, aiming to support users with varying motor or sensory limitations. A considerable body of work has demonstrated that haptic cues can be pivotal in motor learning, spatial orientation, and rhythm synchronization. For example, Xu et al. developed a wearable vibrotactile system that improves postural control through directional feedback, showing encouraging outcomes in both static balance tasks and dynamic movements (Xu et al., 2017). Similarly, Kalff et al. found that gait-synchronized vibrotactile feedback enhances mobility results in subjects with lower-limb amputations, demonstrating how real-time tactile feedback can facilitate neuromotor rehabilitation and entrainment (Kalff et al., 2024).
@@
Beyond medical rehabilitation, haptic feedback has been increasingly integrated into locomotion assistance systems. In the domain of gait training, De Angelis et al. performed an exhaustive analysis and concluded that vibrotactile stimulation not only supports postural stability but also encourages proactive participation by providing non-intrusive cues that users can internalize (De Angelis et al., 2021). Mahmud et al. extended this notion to immersive environments by incorporating vibrotactile signals into virtual reality standing balance training programs, showing that tactile stimuli could effectively substitute or augment visual cues (Mahmud et al., 2022a, b).
@@
Wearable haptic systems have also been used to facilitate socially coordinated behaviors. Lisini et al., for instance, proposed a wearable solution for remote social walking, allowing pairs of persons to preserve synchronized step rhythms even while being geographically separated (Baldi et al., 2020). Their work illustrates the potential of vibrotactile entrainment in interpersonal contexts, suggesting broader applications beyond solo rehabilitation, including shared physical activities like dancing, walking, and running.
@@
However, most prior work focuses on navigation, balance, or motor control in constrained indoor settings, often with the goal of compensating for deficits or preventing falls. Assistive devices designed for the visually impaired, such as Wayband and other commercial systems, have largely targeted directional guidance, providing users with cues for pathfinding or obstacle avoidance through vibrotactile feedback embedded in wearables like shoes or belts. While effective for spatial navigation, these systems are not optimized for interpersonal movement synchronization, particularly in high-tempo and coordination-intensive scenarios like running with a guide (Lind et al., 2020; Islam, 2018).
@@
Moreover, despite the exhibited promise of entrainment theory in social and sensorimotor contexts, few systems have explicitly operationalized this concept into shared, bidirectional pacing tools for co-locomotion. The majority of existing solutions emphasize asymmetrical interaction models, where one user leads and the other passively follows. This dynamic limits autonomy and co-agency, especially for visually impaired individuals who may wish to take a more active role in the running experience (Islam and Lim, 2022).
@@
Our work bridges multiple strands of this literature—rehabilitation, wearable haptics, social entrainment, and assistive technology—to fill a void: the lack of tools that support symmetric, co-tempo running experiences for visually impaired individuals. RunPacer introduces a new category of assistive systems that use synchronized haptic feedback not for navigation, but for enabling mutual rhythm alignment. This repositioning shifts the emphasis from guidance to collaboration, reframing the role of the assistive device as a facilitator of shared agency and embodied communication (Shull and Damian, 2015; Ballardini et al., 2020).
3. System Design
@@
Figure 1. Overview of the RunPacer system, composed of three interconnected modules that operate in tandem to support synchronized co-running.
@@
In either case, a synchronized pulse stream is transmitted to both runners via a low-latency Bluetooth channel. The synchronization system is engineered to operate with under 100 milliseconds of delay, which is essential for preserving the perceptual integrity of the rhythm and maintaining entrainment (Schroeder et al., 2025). The cadence engine is modular and platform-agnostic, supporting integration across smartwatch ecosystems using standardized communication APIs.
@@
Vibrotactile Feedback Delivery, transforms the rhythmic signals into distinct, synchronized pulses using the built-in haptic motors of each smartwatch (Georgiou et al., 2020). These pulses are short, crisp, and calibrated to provide perceptual clarity without becoming intrusive or fatiguing.
The key design parameters include pulse duration, intensity, and waveform configuration. The pulse duration is optimized between 100–200 milliseconds to ensure clarity during motion. Pulse intensity is adjustable to accommodate ambient noise levels, individual preferences, and individual skin sensitivity (Xu et al., 2017). The pulse waveform is configured to deliver single-tap pulses, which were favored over longer or multi-tap patterns in our preliminary user testing.
@@ 7. Conclusion
The contribution of RunPacer lies not only in its specific implementation, but also in the broader design implications it offers for inclusive movement interfaces. It reframes co-running as a bidirectional, rhythmic, and social interaction, leveraging entrainment principles and vibrotactile cues to support mutual adaptation rather than passive following. This reframing has possible implementations beyond running, including co-walking, synchronized rehabilitation, or embodied collaborative learning (Ballardini et al., 2020).